# News is More than a Collection of Facts: Moral Frame Preserving News Summarization

### v1

@@ B.1.1 Summary generation
Table B1 lists the hyperparameters used for the summarization models. If a parameter is not present in the table, the default value supplied by the framework is used. The system content is set to ‘‘You are a news summarizer assistant” for the Plain method and to “You are a news summarizer assistant and a moral expert” for the other four methods.
@@
Table B1: Hyperparameters used for the zero-shot generation.
@@ Appendix C Human evaluations
We provide additional details on the crowd and expert evaluations.
@@ C.1 Crowd evaluation
We provide additional information on the crowd evaluation described in Section 4.4.2. All exact instructions and informed consent forms are provided as supplemental material.
@@
We hosted our evaluation task on the Qualtrics111www.qualtrics.com platform. After obtaining informed consent, we introduced the workers to the concept of moral framing and provided a brief tutorial on the highlighting mechanism by instructing the workers to highlight spans of a mock article with their cursor. Next, they were shown the spans they had highlighted under a mock summary and instructed on the sliding mechanism, to be used to indicate the extent to which the corresponding highlighted morally framed span is preserved in the summary, ranging from 1 (Not Present) to 5 (Clearly Present). Subsequently, they were shown an article followed sequentially by its five summaries. They were asked to highlight all spans of article text they deemed morally framed. For each highlighted span, a slider would be created under each summary, which they would use to indicate the extent to which the corresponding highlighted span of text was preserved in that summary.
@@
18-2425-3435-4445-5455-64$\geq$65002020202040404040CountAge
AsianBlackMixedWhiteOther002020202040404040CountEthnicity
FemaleMaleOther002020202040404040CountGender
@@ C.2 Expert evaluation
The eight experts were contacted through the authors’ network. They are all researchers (including four PhD, three postdoctoral, and one assistant professor) in the age range 26-35 and based in Europe, conducting research in social sciences in the media domain, including behavioral sciences, political communication, and philosophy. We obtained informed consent from all experts (the informed consent form is provided as supplemental material).
@@ D.1 Automated evaluation
We generate summaries with the Plain, Direct, CoT, and Oracle methods for all articles in the dataset, and with Class for the articles in the test set. In Section 5.1 we report the automated evaluation on the test set, while Table D1 shows the results on the full dataset (except for the Class method, for which the evaluation is only performed on the test set). The average MC is 19.42 in the complete dataset and 19.08 in the test set. We observe that the automated evaluation results’ trends in the test set align with those in the full dataset. Thus, since the generation of summaries with the Class method and the human evaluation are performed only on the test set, in the paper we discuss the results on the test set.
@@
Table D1: Results of the automated evaluation of the summaries of the complete dataset.
Table D2 reports the average length of the generated summaries. We observe that the requested maximum summary length (200 words) is generally respected but with variations across methods and models. Command R+ generates consistently longer summaries than Llama-3-70B and DeepSeek R1, in several cases exceeding the word limit. We observe similar trends for Llama-3-70B and DeepSeek R1, with Oracle and Class leading to the longest summaries, likely due to the efforts in preserving the identified moral-laden words. However, despite the instructions to preserve moral-laden words, CoT results in shorter summaries than the other two word-preserving approaches. This is likely due to the number of moral-laden words that the summarizer is tasked to preserve in the summary—the median number of such words is 11 for CoT, 16.5 for Class, and 21.5 for Oracle. Not being tasked with these instructions, Plain and Direct lead to shorter summaries.
Dataset portion Model Plain Direct CoT Oracle Class
Complete dataset Llama-3-70B 153.9 163.0 151.5 175.3 -
Command R+ 172.9 189.6 196.2 193.9 -
DeepSeek R1 146.9 159.1 159.5 172.9 -
Test set Llama-3-70B 153.0 161.4 152.4 174.6 173.9
Command R+ 173.8 191.7 201.3 197.1 194.7
DeepSeek R1 146.8 162.2 159.2 174.3 173.2
Table D2: Average length of the generated summaries.
D.2 Examples
D.2.1 Moral-laden word preservation in the summary

### v2

@@ Appendix A Prompts
We report the complete version of the prompts used to generate summaries.
@@ B.1.1 Summary generation
Table B1 lists the hyperparameters used for the summarization models. If a parameter is not present in the table, the default value supplied by the framework is used. The system prompt is set to “You are a news summarizer assistant” for the Plain method and to “You are a news summarizer assistant and a moral expert” for the other four methods.
@@
Table B1: Hyperparameters used for the zero-shot summarization.
@@ Appendix C Human evaluations
We provide additional details on the crowd and expert evaluations. All exact instructions and informed consent forms are provided as supplemental material (Liscio et al., 2025b).
@@ C.1 Crowd evaluation
We provide additional information on the crowd evaluation described in Section 4.4.2.
@@
We hosted our evaluation task on the Qualtrics222www.qualtrics.com platform. After obtaining informed consent, we introduced the workers to the concept of moral framing and provided a brief tutorial on the highlighting mechanism by instructing the workers to highlight spans of a mock article with their cursor. Next, they were shown the spans they had highlighted under a mock summary and instructed on the sliding mechanism, to be used to indicate the extent to which the corresponding highlighted morally framed span is preserved in the summary, ranging from 1 (Not Present) to 5 (Clearly Present). Subsequently, they were shown an article followed sequentially by its five summaries. They were asked to highlight all spans of article text they deemed morally framed. For each highlighted span, a slider would be created under each summary, which they would use to indicate the extent to which the corresponding highlighted span of text was preserved in that summary.
@@
[…]
AsianBlackMixedWhiteOther020204040CountEthnicity
FemaleMaleOther020204040CountGender
@@ C.2 Expert evaluation
The eight experts were contacted through the authors’ network. They are all researchers (including four PhD, three postdoctoral researchers, and one assistant professor) in the age range 26-35 and based in Europe, conducting research in social sciences in the media domain, including behavioral sciences, political communication, and philosophy. We obtained informed consent from all experts.
@@ D.1 Automated evaluation
We generate summaries with the Plain, Direct, CoT, and Oracle methods for all articles in the dataset, and with Class for the articles in the test set. In Section 5.1 we report the automated evaluation on the test set, while Table D1 shows the results on the full dataset (except for the Class method, for which the evaluation is only performed on the test set). The average MoralCount is 19.42 in the complete dataset and 19.08 in the test set. The average article length in the complete dataset is 719.73. We observe that the automated evaluation results’ trends in the test set align with those in the full dataset. Thus, since the generation of summaries with the Class method and the human evaluation are performed only on the test set, in the paper we discuss the results on the test set.
@@
Summary Length Llama-3-70B 153.9 163.0 151.5 175.3
Command R+ 172.9 189.6 196.2 193.9
DeepSeek R1 146.9 159.1 159.5 172.9
Table D1: Results of the automated evaluation of the summaries of the complete dataset. As detailed in Section 4.4.1, we report three reference-free automated evaluation metrics, two novel metrics to measure the preservation of moral-laden words, and the average length of the summaries. $\uparrow$ indicates that higher scores are better, $\downarrow$ indicates the opposite.
D.2 Inter-annotator agreement
As noted in Section 5.2, traditional inter-annotator agreement metrics are not appropriate in this context, as annotators highlighted and rated different text spans. Instead, we offer a proxy measure to approximate agreement. To illustrate this, we walk through a hypothetical annotation example that clarifies the process in detail. Table D2 shows a fictitious example of the crowd annotations for the summaries generated through one of the five methods.
Art. ID Ann. ID Likert scores Mean ann. score Mean art. score SD art. score
0 a [2,2,3,5] 3 3.17 0.17
0 b [3,3,4] 3.33
1 c [3,4,5] 4 3.6 0.4
1 d [1,3,5,4,3] 3.2
$\vdots$ $\vdots$ $\vdots$ $\vdots$
59 x [2,5] 3.5 2.95 0.55
59 y [3,3,3,1,2] 2.4
Table D2: Fictitious example of the crowd annotations for the 60 test set summaries generated through one of the five prompting methods.
In this example, two annotators (a and b) annotated article 0. Each of the two annotators highlighted a different number of morally framed spans in the article (4 and 3, respectively) and thus assigned different numbers of Likert scores. Calculating the inter-annotator agreement of these scores is not possible, as these scores were assigned to different items (i.e., different text spans). We then average these Likert scores, resulting in one score per annotator (3 and 3.33, respectively). Then, we average the two annotators’ scores into one article score (3.17). We perform this for all 60 articles in the test set, and report in Table 4 the mean and standard deviation of these 60 scores (in this fictitious case, we would report the mean and standard deviation of the 60-dimensional vector [3.17, 3.6, …, 2.95]).
Additionally, we can calculate the standard deviation between the two mean annotator scores assigned to one article (in the case of article 0 in this example, the standard deviation of the vector [3, 3.33], which is 0.17). We can then compute the mean and standard deviation of the standard deviation vector (in the fictitious example above, of the vector [0.17, 0.4,…,0.55]).
Method Mean SD
Plain 0.80 $\pm$ 0.53
Direct 0.59 $\pm$ 0.47
CoT 0.59 $\pm$ 0.46
Oracle 0.55 $\pm$ 0.45
Class 0.57 $\pm$ 0.48
Table D3: Mean and SD of the SD vector of the results.
Table D3 reports the mean and standard deviation of the standard deviation vector of our experimental results across the five methods. These results show that the Plain approach leads to larger differences in average Likert scores among annotators when compared to the other approaches, which are comparable to each other. Furthermore, the magnitude of the mean standard deviation appears relatively low (although difficult to judge as no baseline is available), hinting at a consistency among annotators.
D.3 Examples
We provide additional examples of annotations in the EMONA dataset, misclassification of moral-laden words, and motivations provided by the experts in support of their evaluations.
D.3.1 Annotation examples
Table D5 displays two additional examples of event annotations in the EMONA dataset, adapted from the original paper (Lei et al., 2024), similar to the example shown in Figure 1. Table D5 lists the ten MFT moral foundations used to annotate moral-laden events.
Annotation examples
In a succinct speech [non-moral], the new president Donald Trump told [non-moral] Americans: “The time for empty talk [cheating] is over. Now arrives the hour of action [authority].”
Texas Gov. Rick Perry is on the attack [non-moral], claiming [non-moral] in a TV ad that President Obama is waging a war [degradation] on religion, and he is the GOP candidate that can defend [purity] faith in America.
Table D4: Example annotations in the EMONA dataset. In brackets, non-moral, virtue, and vice annotations.
Foundation Definition
Care/
Harm
Support for care for others/
Refrain from harming others
Fairness/
Cheating
Support for fairness and equality/
Refrain from cheating or exploiting others
Loyalty/
Betrayal
Support for prioritizing one’s inner circle/
Refrain from betraying the inner circle
Authority/
Subversion
Support for respecting authority and
tradition/
Refrain from subverting authority or
tradition
Purity/
Degradation
Support for the purity of sacred entities/
Refrain from corrupting such entities
Table D5: The MFT moral foundations (virtue/vice).
D.3.2 Moral-laden word preservation in the summary